2 A 50-kg person was moving on a 20-kg movable platform on top of a smooth surface such as ice. A force of 70-N was applied to the platform and the platform responded instantly at 4 m/s/s. 2a Find the minimum force (magnitude and direction) that the person exerted on the platform. 2b Determine the dancer moving direction, right or left, as the platform responded instantly at 4m/s/s. Explain.
Answer #1

2a) the mini mum iurex: 1% , F+Ff=ma 70=(20)(4) F-80-70 = 1 ON The direction is right 2b) since the plate form start moving 4m/s is right Therefore, the denser direction is left